# Quadratic curvature corrections to 5-dimensional Kerr-AdS black hole thermodynamics by background subtraction method

### v2

@@ I Introduction
In the Euclidean approach to black hole thermodynamics, there have been two methods proposed to obtain the finite Gibbs free energy of stationary black holes and the relevant thermodynamic quantities. One is the background subtraction method invented by Hawking and his companions with the inception of the Euclidean approach to quantum gravity[1, 2, 3, 4, 5], and the other is the covariant counterterm method, also dubbed as holographic renormalization method because it is motivated by Anti-de Sitter/Conformal Field Theory (AdS/CFT) correspondence[6, 7, 8]. Both methods are believed to have their advantages and disadvantages, respectively. Although the holographic renormalization method is supposed to be universally applicable, the involved calculation is cubersome because it needs complicated dimension-dependent boundary terms in addition to the generalized Gibbons-Hawking-York term for a generic diffeomorphism covariant gravity theory. On the other hand, the background subtraction method is much simpler, but it is suspected to have a rather restrictive applicability in the sense that the induced boundary metric of the reference spacetime cannot be the exactly the same as that of stationary black holes. Gratefully, a very progress has been made most recently regarding the applicability of the background subtraction method in [9], where it is found that the criterion for the applicability of the background subtraction method does not require the aforementioned induced boundary metrics be exactly the same, instead the background subtraction method turns out to be as applicable as the covariant counterterm method indeed. However, such an applicability is explicitly demonstrated only for the 4-dimensional rotating black hole. The purpose of this paper is two-fold. For one thing, we shall show that the background subtraction method is also applicable to the 5-dimensional Kerr-AdS black hole for the Einstein’s gravity and its higher derivative corrections. For another, we shall apply the background subtraction method to calculate the quadratic curvature corrections to the 5-dimensional Kerr-AdS black hole thermodynamics. Among the recent efforts in the investigation of the higher derivative corrections to black hole thermodynamics [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], the authors in [23] has already obtained such quadratic curvature corrections to the Gibbs free energy for the 5-dimensional Kerr-AdS black hole. But the method they use is the holographic renormalization method. So this provides us with a very opportunity to compare our result with theirs. The resulting exact agreement between them further substantiates the applicability of the background subtraction method.
@@ II Criterion for the applicability of background subtraction method in $F(R_{abcd})$ gravity
As shown in [9] for the Killing vector field $\xi$ normal to the stationary black hole horizon and tangential to $\Gamma$ , the first law of black hole thermodynamics
$T\delta S=\delta[H_{\xi}]$ (9)
is satisfied if and only if
$\int_{\mathcal{S}}\xi\cdot[\mathbf{F}]=0\quad\texttt{as}\quad\Gamma\rightarrow\infty$ (10)
@@
The condition (10) together with the finiteness condition for $H_{\xi}]$ as $\Gamma\rightarrow\infty$ constitutes the necessary and sufficient condition for the applicability of the background subtraction method. If such a criterion is satisfied, then the Gibbs free energy of the stationary black holes can be expressed as
@@ III Applicability of background subtraction method to 5-dimensional asymptotically AdS spacetimes
where $T_{bc}=-\frac{1}{8\pi}(K_{bc}-Kh_{bc})$ is the well-known Brown-York boundary energy-momentum tensor. On the other hand, the above Lagrangian form allows the following Kerr-AdS black hole solution
@@ Acknowledgements.
This work is partially supported by the National Key Research and Development Program of China with Grant No. 2021YFC2203001 as well as the National Natural Science Foundation of China with Grant No. 12361141825 and 12475049.

### v3

@@ I Introduction
In the Euclidean approach to black hole thermodynamics, there have been two methods proposed to obtain the finite Gibbs free energy of stationary black holes and the relevant thermodynamic quantities. One is the background subtraction method invented by Hawking and his companions with the inception of the Euclidean approach to quantum gravity [1, 2, 3, 4, 5, 6], which has been used to successfully extract various higher derivative corrections to the static black hole thermodynamics [7, 8, 9, 10, 11, 12, 13, 14, 15]. The other is the covariant counterterm method, also dubbed as holographic renormalization method because it is motivated by Anti-de Sitter/Conformal Field Theory (AdS/CFT) correspondence [16, 17, 18]. Both methods are believed to have their advantages and disadvantages, respectively. Although the holographic renormalization method is supposed to be universally applicable, the involved calculation is cubersome because it needs complicated dimension-dependent boundary terms in addition to the generalized Gibbons-Hawking-York term for a generic diffeomorphism covariant gravity theory. Compared to this, the background subtraction method is much simpler, but it is suspected to have a rather restrictive applicability in the sense that the induced boundary metric of the reference spacetime cannot be the exactly the same as that of stationary black holes. Gratefully, one very progress has been made most recently regarding the applicability of the background subtraction method in [19], where it is found that the criterion for the applicability of the background subtraction method does not require the aforementioned induced boundary metrics be exactly the same and the background subtraction method turns out to be as applicable as the covariant counterterm method indeed. However, such an applicability is explicitly demonstrated only for the 4-dimensional rotating black hole. The purpose of this paper is two-fold. For one thing, we shall show that the background subtraction method is also applicable to the 5-dimensional Kerr-AdS black hole for the Einstein’s gravity and its higher derivative corrections. For another, we shall apply the background subtraction method to calculate the quadratic curvature corrections to the 5-dimensional Kerr-AdS black hole thermodynamics. Among the recent efforts in the investigation of the higher derivative corrections to black hole thermodynamics [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38], the authors in [33] have already obtained such quadratic curvature corrections to the Gibbs free energy for the 5-dimensional Kerr-AdS black hole. But the method they use is the holographic renormalization method. So this provides us with a very opportunity to compare our result with theirs. The resulting exact agreement between them further substantiates the applicability of the background subtraction method.
@@ II Criterion for the applicability of background subtraction method in $F(R_{abcd})$ gravity
As shown in [19] for the Killing vector field $\xi$ normal to the stationary black hole horizon and tangential to $\Gamma$, we have
$T\delta S=\delta[H_{\xi}]+\int_{\mathcal{S}}\xi\cdot[\mathbf{F}]$ (9)
@@
Note that unlike the first term on the right handed side of Eq. (9), the second term is not a complete variation, So to have our familiar first law of black hole thermodynamics, we require
$\int_{\mathcal{S}}\xi\cdot[\mathbf{F}]=0\quad\texttt{as}\quad\Gamma\rightarrow\infty.$ (14)
This requirement together with the finiteness condition for $H_{\xi}]$ as $\Gamma\rightarrow\infty$ constitutes the necessary and sufficient condition for the applicability of the background subtraction method. If such a criterion is satisfied, then the Gibbs free energy of the stationary black holes can be expressed as
@@ III Applicability of background subtraction method to 5-dimensional asymptotically AdS spacetimes
where $T_{bc}=-\frac{1}{8\pi}(K_{bc}-Kh_{bc})$ is the well-known Brown-York boundary energy-momentum tensor. On the other hand, the above Lagrangian form allows the rotating Kerr-AdS black hole solution, which was firstly obtained in [5]. For our purpose, we like to write this solution as follows [6]
@@ Acknowledgements.
This work is partially supported by the National Key Research and Development Program of China with Grant No. 2021YFC2203001 as well as the National Natural Science Foundation of China with Grant No. 12361141825, 12475049, and 12575047.